# Aoxiang MA    Salah GHAMIZI    Jun CAO Pedro RODRIGUEZ CORTES

### v1

@@ I Introduction
Three-phase unbalanced distribution grids present unique challenges for real-time battery optimal operational dispatch. Unlike transmission systems characterized by balanced three-phase conditions, distribution grids exhibit significant phase imbalances due to uneven load distribution across phases, creating substantial phase load differences [sun2023analysis]. These imbalances create significant phase-to-phase coupling effects via mutual impedances that single-phase approximations fail to capture [kersting2010distribution]. The integration of distributed energy resources and battery energy storage systems further complicates the operation, requiring both accurate multi-variable prediction and constraint-aware dispatch strategies [gangwar2024energy].
Traditional approaches rely on three-phase OPF solvers combined with a Multi-step optimizer for battery optimal dispatch [opends]. Although these methods are rigorous and physically sound, they require long computation time per optimization cycle, fundamentally limiting real-time applications in dynamic grid environments. Graph neural networks (GNNs) have emerged as scalable alternatives, exploiting grid topology through localized parameterization where parameters are defined on nodes and edges rather than globally [owerko2020optimal]. This localized approach enables superior performance for predicting voltage magnitudes, phase angles, and power injections compared to fully connected architectures [donon2020neural]. Recent advances have extended GNNs to multi-task frameworks, simultaneously forecasting multiple power system variables while capturing spatial dependencies in complex grid topologies [hansen2021power].
Heterogeneous GNNs have advanced constraint-aware control capabilities. Ghamizi et al. [ghamizi2024opf] proposed OPF-HGNN with type-specific embeddings and differentiable constraints, achieving 98% constraint satisfaction on balanced networks with two orders of magnitude speedup. Ma et al. [ma2025hil] validated real-time operation through hardware-in-the-loop (HIL) simulations using the SafePowerGraph-HIL framework, demonstrating practical feasibility for grid deployment.
However, existing HIL-tested heterogeneous GNN approaches remain limited to single-phase or balanced-condition scenarios [ma2025hil], lacking explicit battery constraint enforcement in unbalanced systems where SoC and C-rate violations present critical operational risks. While PowerFlowMultiNet [ghamizi2024multigraph] addresses unbalanced three-phase power flow prediction using multigraph representations, and physics-informed approaches [qin2025physics] enable dynamic voltage regulation, these works focus primarily on power flow prediction and voltage control rather than constraint-aware battery dispatch in unbalanced distribution grids.
@@ II-B1 Graph Embedding
We model lines as nodes, given the effectiveness and superiority of this representation [ghamizi2024multigraph]. This heterogeneous approach enables the GNN to learn how line impedances and mutual coupling terms affect voltage propagation across phases.
@@ III Empirical Study
We use PandaPower[thurner_pandapower_2018] for topology validation, 3-phase optimal power flow (OPF) ground truth generation. The performance study uses 8,000 training datasets and 2,000 validation datasets.

### v2

@@ I Introduction
Three-phase unbalanced distribution grids present unique challenges for real-time battery optimal operational dispatch. Unlike transmission systems characterized by balanced three-phase conditions, distribution grids exhibit significant phase imbalances due to uneven load distribution across phases, creating substantial phase load differences [11]. These imbalances create significant phase-to-phase coupling effects via mutual impedances that single-phase approximations fail to capture [6]. The integration of distributed energy resources and battery energy storage systems further complicates the operation, requiring both accurate multi-variable prediction and constraint-aware dispatch strategies [2].
Traditional approaches rely on three-phase OPF solvers combined with a Multi-step optimizer for battery optimal dispatch [10]. Although these methods are rigorous and physically sound, they require long computation time per optimization cycle, fundamentally limiting real-time applications in dynamic grid environments. Graph neural networks (GNNs) have emerged as scalable alternatives, exploiting grid topology through localized parameterization where parameters are defined on nodes and edges rather than globally [8]. This localized approach enables superior performance for predicting voltage magnitudes, phase angles, and power injections compared to fully connected architectures [1]. Recent advances have extended GNNs to multi-task frameworks, simultaneously forecasting multiple power system variables while capturing spatial dependencies in complex grid topologies [5].
Heterogeneous GNNs have advanced constraint-aware control capabilities. Ghamizi et al. [4] proposed OPF-HGNN with type-specific embeddings and differentiable constraints, achieving 98% constraint satisfaction on balanced networks with two orders of magnitude speedup. Ma et al. [7] validated real-time operation through hardware-in-the-loop (HIL) simulations using the SafePowerGraph-HIL framework, demonstrating practical feasibility for grid deployment.
However, existing HIL-tested heterogeneous GNN approaches remain limited to single-phase or balanced-condition scenarios [7], lacking explicit battery constraint enforcement in unbalanced systems where SoC and C-rate violations present critical operational risks. While PowerFlowMultiNet [3] addresses unbalanced three-phase power flow prediction using multigraph representations, and physics-informed approaches [9] enable dynamic voltage regulation, these works focus primarily on power flow prediction and voltage control rather than constraint-aware battery dispatch in unbalanced distribution grids.
@@ II-B1 Graph Embedding
We model lines as nodes, given the effectiveness and superiority of this representation [3]. This heterogeneous approach enables the GNN to learn how line impedances and mutual coupling terms affect voltage propagation across phases.
@@ III Empirical Study
We use PandaPower[12] for topology validation, 3-phase optimal power flow (OPF) ground truth generation. The performance study uses 8,000 training datasets and 2,000 validation datasets.